Find an equation of the line that is tangent to the graph of f and parallel to the given line.

function: f(x) = x2 − 6
line: 2x + y = 0
Answer #1

Solve for y:

2x+y=0


Add -2x to both sides

y=-2x


The slope is m=-2


Find the derivative for f(x) = x2 − 6

f'(x)=2x2-1-0=2x


Make f'(x)=m

2x=-2


Divide by 2 both sides

x=-1


Substitute in f(x)

f(-1)=(-1)2-6=1-6=-5


Use the formula

y=m(x-x0)+f(a)


y=-2(x-(-1))+(-5)=-2(x+1)-5


Simplify

y=-2x-2-5


y=-2x-7


image.png

answered by: robertoxyz20